IF X=Y-1 AND Y=4-2X; SOLVING BY THE SUBSTUTION METHOD; WHAT ARE THE VALUES FOR FOR X AND Y?
Answer by SoOoDumb(14)

Get both equations to y-intercept form. (Optional) y=mx+b
x=y-1
x-y=-1
-y=-x-1(-1)
y=x+1
y=4-2x
y=-2x+4

Now that x+1= y, and -2x+4 =y, we can say that x+1=-2x+4.
x+1=-2x+4
x=-2x+3
3x=3
x=1
Now that we have the x, we can find y.
y=x+1
y=(1)+1
y=2
Solution: (1,2)